OpenStudy (anonymous):

Identify the initial amount a and the growth factor b in the exponential function. f(x) = 410 * 4.9^x

OpenStudy (anonymous):

HELP

OpenStudy (zzr0ck3r):

\(f(x) = a*b^x\) \(a\) is the initial amount and \(b\) is the growth factor

OpenStudy (zzr0ck3r):

So what is the answer?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

idk how bout u tel me lol

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

ANYONE

OpenStudy (zzr0ck3r):

read the rules

OpenStudy (anonymous):

omg nick zzsomething practically gave u the answer. look at the equation and look at mr. z's guided equation. what is b?

OpenStudy (anonymous):

ok??

OpenStudy (zzr0ck3r):

it cant be any easier than this you have \(f(x) = 410 * 4.9^x\) i shoed you how to do it for \(f(x) = a*b^x\)

OpenStudy (zzr0ck3r):

we are not here to give answers, we are here to help.

OpenStudy (zzr0ck3r):

it would be strange if I had a desire to do your homework for you:)

OpenStudy (anonymous):

but i dont understand

OpenStudy (zzr0ck3r):

dude then you are not trying. A first grader could do this \(f(x) = 410 * 4.9^x\) \(f(x) = a*b^x\) what the a in your equation?

OpenStudy (anonymous):

ANWERS- A. 410, 4.9 B. 4.9, 410

OpenStudy (anonymous):

so its A

OpenStudy (zzr0ck3r):

we arnt going to tell you the answer

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

thx Mr. Dude

OpenStudy (anonymous):

**** nick. if u have eyes then u would know